OpenStudy (anonymous):

solve y=3x-11 and 2x+3y=0 using substitution method? please quick answers and show work!!!

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

\(\Large 2x+3y=0 \Leftrightarrow 2x+3(3x-11)=0 \Leftrightarrow 11x-33=0 \Leftrightarrow x=\frac{33}{11} =3 \)

OpenStudy (anonymous):

\(\Large x=3\ and\ y= 9-11=-2\)